OpenStudy (anonymous):

John earns $13.00 each week from his paper route.He earns $.15 for each daily paper and $.35 for each sunday paper he delivers.He delivers twice as many as daily papers as sunday papers. How many of each type does he delivers each week.

OpenStudy (ajprincess):

Let the number of daily papers be 2x. then the number of sunday papers is x. 2x*0.15+x*0.35=13 0.30x+0.35x=13 0.65x=13 65x/100=13 65x=13*100 65x=1300 x=1300/65 x=20

OpenStudy (ajprincess):

number of daily papers =2*20=40 number of sunday papers =20 Is it clear?

OpenStudy (anonymous):

I'm yet to double check. Am having lunch now. Are you going to be on for a while?

OpenStudy (ajprincess):

ya will be here for another 15 minutes.

OpenStudy (anonymous):

Thanks. But am in the USA and I don't know the time difference with where you are. It may take me a while to get to this.